# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 4.1. Preliminaries
We recall some classical definitions from birational geometry and the Minimal Model Program using the terminology of [undefap, Section 2].
@@
Let $f:X\to S$ be a fibration between $\mathbb{Q}$-factorial normal varieties. Two divisors $D$ and $D^{\prime}$ on $X$ are said to be $f$-linearly equivalent if their difference is linearly equivalent to the pullback of a divisor on $S$. We denote by $\mathop{\rm Pic}\nolimits(X/S)$ the relative Picard group (of $f$), and for $\mathbb{K}=\mathbb{Q}$ or $\mathbb{R}$, we set
@@
Let $f:X\to S$ be a fibration between $\mathbb{Q}$-factorial normal varieties. Two $\mathbb{R}$-divisors $D$ and $D^{\prime}$ on $X$ are said to be $f$-numerically equivalent if $D\cdot C=D^{\prime}\cdot C$ for every curve $C\subset X$ such that $f(C)$ is a point. We denote by $N^{1}(X/S)$ the relative Néron-Severi group (of $f$), and for $\mathbb{K}=\mathbb{Q}$ or $\mathbb{R}$, we set
@@
We will say that a minimal model $(Y,\Delta_{Y})$ over $S$ is a good minimal model if, in addition, $K_{Y}+\Delta_{Y}$ is semiample over $S$.
In the next section, we will also need to refer to the weak cone conjecture; we therefore take the opportunity in this preliminary section to introduce its definition following [undefaq] and [undefau].
@@
Let $V$ be a real vector space, $C\subset V$ a cone and $\rho:\Gamma\hookrightarrow\mathop{\rm GL}\nolimits(V)$ an injective group homomorphism. Let $\Pi\subset C$ be a (rational) polyhedral cone. Suppose that $\Gamma$ acts on $C$. Then $\Pi$ is called a weak (rational) polyhedral fundamental domain for $C$ for the action of $\Gamma$ if
@@
  • $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$ has a weak rational polyhedral fundamental domain for the action of $\mathop{\rm PsAut}\nolimits(X/S,\Delta)$.
  • $\mathop{\rm Nef}\nolimits(X/S)^{+}$ has a weak rational polyhedral fundamental domain for the action of $\mathop{\rm Aut}\nolimits(X/S,\Delta)$.
@@ 4.2. Proof of the relative cone conjecture
We first recall the following lemma which relies on [undefau, Lemma 3.7] and the proof of [undefau, Theorem 1.3].
Lemma 4.11 (Lemma 5.1 - [undefai]).
@@
Let $f:(X,\Delta)\to S$ be a $K$-trivial fiber space $f:(X,\Delta)\to S$ (cf. Definition 1.1) whose very general fibre $X_{s}$ has a vanishing irregularity $q(X_{s})$, and assume the base $S$ is $\mathbb{Q}$-factorial. As it was observed in [undefai, Remark 1.6] and in the remark following [undefai, Lemma 5.1], the following properties hold:
@@
Let $X$ and $S$ be complex quasi-projective varieties, and let $f:X\to S$ be a projective fibration. Denote by $\eta\in S$ the generic point of the base and set $F:=\mathbb{C}(S)$. If the very general fibre of $f$ is a complex projective primitive symplectic variety with $\mathbb{Q}$-factorial terminal singularities, then the generic fibre $X_{\eta}$ is a projective primitive symplectic variety over $F$ with $\mathbb{Q}$-factorial terminal singularities.
@@
Let $\overline{F}$ be an algebraic closure of $F$, and let $\overline{\eta}\in S$ be the image of the composition $\mathop{\rm Spec}\nolimits(\overline{F})\to\mathop{\rm Spec}\nolimits(F)\to S$. We call $\overline{\eta}$ the geometric generic point of $S$ and $X_{\overline{\eta}}$ the geometric generic fibre of $f$. By Lemma 2.4, in order to show that $X_{\eta}$ is a projective primitive symplectic variety over $F$, it suffices to prove that $X_{\overline{\eta}}$ is a primitive symplectic variety over $\overline{F}$.
One way to proceed is to exploit the fact that we are working over $\mathbb{C}$ and to apply [undefaax, Lemma 2.1]. We can then find a closed point $s\in S$ such that $X_{s}$ is a complex primitive symplectic as in the statement of this lemma, an (abstract) field isomorphism $\phi:\mathbb{C}\to\overline{F}$, and an isomorphism of schemes $\alpha:X_{\overline{\eta}}\to X_{s}$ over $\mathop{\rm Spec}\nolimits(\phi)$. It follows that
@@
is a projective primitive symplectic variety over $\overline{F}$ by Lemma 2.3. Moreover, the isomorphism $\alpha$ induces isomorphisms between the divisor class groups and the Picard groups of $X_{s}$ and $X_{\overline{\eta}}$, and the following diagram is commutative.
@@
It follows that $X_{s}$ is $\mathbb{Q}$-factorial if and only if $X_{\overline{\eta}}$ is $\mathbb{Q}$-factorial. Since $K_{X_{\overline{\eta}}}=\alpha^{*}K_{X_{s}}$ we also see that the singularities of $X_{s}$ are terminal if and only if those of $X_{\overline{\eta}}$ are also terminal. Therefore, by [undefq, Lemma 6.18] and [undefao, Proposition 2.15], the singularities of $X_{\eta}$ are $\mathbb{Q}$-factorial and terminal, and that completes the proof. ∎
We now state a weaker version of [undefs, Lemma 2.17] in the singular setting, which will be useful in what follows.
@@
Since the interiors of the cones are the same, we only have to check that the boundary of the left-hand cone is included in the right-hand cone. Let $\alpha\in\partial\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$ and, without loss of generality, suppose that $\alpha$ is integral. Since $\overline{\mathop{\rm BAmp}\nolimits(X)}=\mathop{\overline{\rm Mov}}\nolimits(X)$ by [undefat, Proposition 5.8], there exists a projective primitive symplectic variety $X^{\prime}$ with $\mathbb{Q}$-factorial and terminal singularities, and a birational map $\phi:X\dashrightarrow X^{\prime}$ such that $\phi_{*}\alpha$ is integral and nef. Thanks to our assumptions, $\phi_{*}\alpha$ is semi-ample on $X^{\prime}$, and is therefore effective. Since $\phi$ is an isomorphism in codimension one, $\alpha$ is effective and this concludes the proof. ∎
The following result is an adaptation of [undefai, Proposition 5.5] to the singular setting.
@@
Let $f:X\to S$ be a projective fibration between $\mathbb{Q}$-factorial varieties, and let $X_{s}$ denote a very general fibre of $f$. Let $\eta\in S$ be the generic point of the base, and denote by $j:X_{\eta}\to X$ the inclusion. Then the following assertions hold.
@@
is ample on $X_{\eta}$. Moreover, by [undefae, Corollaire 9.6.4], for each such integer $m$, there exists an open subset $U_{m}\subseteq S$ such that $\left(D+m^{-1}H\right)|_{X_{s}}$ is ample on $X_{s}$ for all $s\in U_{m}$. By (4.1), these open subsets are non-empty. Set $U:=\cap_{m\geq 1}U_{m}$. By construction, $D|_{X_{s}}$ is nef for all $s\in U$, and our hypothesis implies that $D|_{X_{s}}$ is also effective for all $s\in U$. Hence $D\in\mathop{\rm Eff}\nolimits(X/S)$, and therefore $D_{\eta}\in\mathop{\rm Eff}\nolimits(X_{\eta})$.
(2) We first prove the inclusion $\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})\subset\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$. Let $D_{\eta}\in\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})$. By Lemma 4.11, there exists $D\in\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)$ such that $D|_{X_{\eta}}=D_{\eta}$. By [undefai, Lemma 3.4], we have $D\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$, thus $D_{\eta}\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$. For the reverse inclusion $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})\subset\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})$, let $D_{\eta}\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$ and assume, without loss of generality, that $D_{\eta}$ is rational. By Lemma 4.11, we may choose $D\in N^{1}(X/S)_{\mathbb{Q}}$ such that $D|_{X_{\eta}}=D_{\eta}$. Let $H$ be an $f$-ample divisor on $X$. For any integer $m>0$, the divisor
@@
is big on $X_{\eta}$, thus we may write $(D+m^{-1}H)|_{X_{\eta}}=A_{m,\eta}+E_{m,\eta}$ with $A_{m,\eta}\in\mathop{\rm Amp}\nolimits(X_{\eta})$ and $E_{m,\eta}\in\mathop{\rm Eff}\nolimits(X_{\eta})$. Choose rational classes $A_{m}\in N^{1}(X/S)_{\mathbb{Q}}$ and $E_{m}\in\mathop{\rm Eff}\nolimits(X/S)$ such that $(A_{m})|_{X_{\eta}}=A_{m,\eta}$ and $(E_{m})|_{X_{\eta}}=E_{m,\eta}$. As above, for each integer $m>0$, there exists a non-empty open subset $U_{m}\subseteq S$ such that $(D+m^{-1}H)|_{X_{s}}=(A_{m}+E_{m})|_{X_{s}}$ is big for all $s\in U_{m}$. Setting again $U:=\cap_{m\geq 1}U_{m}$, it follows that $D|_{X_{s}}$ is pseudoeffective for all $s\in U$. By [undefas, Theorem 3.2], we may write $D|_{X_{s}}=P_{s}+N_{s}$, with $P_{s}\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{s})$ and $N_{s}\in\mathop{\rm Eff}\nolimits(X_{s})$. Since $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{s})\subseteq\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{s})$ by Lemma 4.14, it follows that $D|_{X_{s}}$ is effective. Therefore, $D$ is $f$-effective, and hence $D_{\eta}$ is effective.
As for the equality $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)=\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)$, the inclusion $\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)\subset\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$ can be proved in the same way as above. The reverse inclusion follows from [undefai, Lemma 3.4], which concludes the proof. ∎
We now have all the ingredients needed to prove our final theorem.
@@
By Lemma 4.13, the generic fibre $X_{\eta}$ is a projective primitive symplectic variety over $F=\mathbb{C}(S)$ with $b_{2}(X_{\eta})\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Theorem A.(2) therefore implies that $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$ admits a rational polyhedral fundamental domain for the action of $\mathop{\rm Bir}\nolimits(X_{\eta})$. Note that, under our assumptions, we have $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})=\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})$ by Proposition 4.15. As observed in [undefai], the proof of [undefau, Theorem 6.1] shows that the weak cone conjecture 4.10 for $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$ implies the weak cone conjecture for $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$. Finally, since we also have $\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)=\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$ by Proposition 4.15, it follows from [undefai, Lemma 2.18] that the relative movable cone conjecture holds for $f:X\to S$.
Finally, by [undefai, Proposition 4.3], there are only finitely many small $\mathbb{Q}$-factorial modifications of $X$ over $S$ up to isomorphism over the base, and the relative nef cone conjecture for each of them. This completes this proof. ∎

### v2

@@ 4.1. Preliminaries
We recall some classical definitions from birational geometry and the Minimal Model Program using the terminology of [undefao, Section 2].
@@
Let $f:X\to S$ be a fibration between $\mathbb{Q}$-factorial normal varieties. Two divisors $D$ and $D^{\prime}$ on $X$ are $f$-linearly equivalent, written $D\sim_{f}D^{\prime}$ or $D\sim_{S}D^{\prime}$, if their difference is linearly equivalent to the pullback of a divisor on $S$. We denote by $\mathop{\rm Pic}\nolimits(X/S)$ the relative Picard group (of $f$), and for $\mathbb{K}=\mathbb{Q}$ or $\mathbb{R}$, we set
@@
Let $f:X\to S$ be a fibration between $\mathbb{Q}$-factorial normal varieties. Two $\mathbb{R}$-divisors $D$ and $D^{\prime}$ on $X$ are $f$-numerically equivalent, written $D\equiv_{f}D^{\prime}$ or $D\equiv_{S}D^{\prime}$, if $D\cdot C=D^{\prime}\cdot C$ for every curve $C\subset X$ such that $f(C)$ is a point. We denote by $N^{1}(X/S)$ the relative Néron-Severi group (of $f$), and for $\mathbb{K}=\mathbb{Q}$ or $\mathbb{R}$, we set
@@
We will say that a minimal model $(Y,\Delta_{Y})$ over $S$ is a good minimal model if, in addition, $K_{Y}+\Delta_{Y}$ is semiample over $S$. A minimal model $(Y,\Delta_{Y})$ together with the birational contraction $\phi$ is called a marked minimal model of $(X,\Delta)$ over $S$.
In the next section, we will also need to refer to the weak cone conjecture; we therefore take the opportunity in this preliminary section to introduce its definition following [undefap] and [undefat].
@@
Let $V$ be a real vector space, $C\subset V$ a cone and $\rho:\Gamma\hookrightarrow\mathop{\rm GL}\nolimits(V)$ an injective group homomorphism. Let $\Pi\subset C$ be a (rational) polyhedral cone. Suppose that $\Gamma$ acts on $C$. Then $\Pi$ is called a weak (rational) polyhedral fundamental domain for $C$ under the action of $\Gamma$ if
@@
  • $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$ has a weak rational polyhedral fundamental domain under the action of $\mathop{\rm PsAut}\nolimits(X/S,\Delta)$.
  • $\mathop{\rm Nef}\nolimits(X/S)^{+}$ has a weak rational polyhedral fundamental domain under the action of $\mathop{\rm Aut}\nolimits(X/S,\Delta)$.
@@ 4.2. Proof of the relative cone conjecture
We first recall the following lemma which relies on [undefat, Lemma 3.7] and the proof of [undefat, Theorem 1.3].
Lemma 4.11 (Lemma 5.1 - [undefah]).
@@
Let $f:(X,\Delta)\to S$ be a $K$-trivial fiber space $f:(X,\Delta)\to S$ (cf. Definition 1.1) whose very general fibre $X_{s}$ has a vanishing irregularity $q(X_{s})$, and assume the base $S$ is $\mathbb{Q}$-factorial. As it was observed in [undefah, Remark 1.6] and in the remark following [undefah, Lemma 5.1], the following properties hold:
@@
Definition 4.13.
Let $f:X\to S$ be a fibration between complex quasi-projective varieties. If $\eta\in S$ denotes the generic point of $S$, we denote by $\overline{\eta}\in S$ the image of the composition $\mathop{\rm Spec}\nolimits\overline{\mathbb{C}(S)}\to\mathop{\rm Spec}\nolimits\mathbb{C}(S)\to S$, where $\mathbb{C}(S)$ is the function field of $S$. We call $\overline{\eta}$ the geometric generic point of $S$, and $X_{\overline{\eta}}$ the geometric generic fibre of $f$.
@@
Let $X$ and $S$ be complex quasi-projective varieties, and let $f:X\to S$ be a projective fibration. Denote by $\eta\in S$ the generic point of $S$ and set $F:=\mathbb{C}(S)$. If the very general fibre $X_{s}$ of $f$ is a complex projective primitive symplectic variety with $\mathbb{Q}$-factorial terminal singularities and $b_{2}(X_{s})\geq 5$, then the generic fibre $X_{\eta}$ (respectively, the geometric generic fibre $X_{\overline{\eta}}$) is a projective primitive symplectic variety over $F$ (respectively, over $\overline{F}$) with $\mathbb{Q}$-factorial terminal singularities and $b_{2}(X_{\eta})\geq 5$.
@@
By Lemma 2.4, in order to show that $X_{\eta}$ is a projective primitive symplectic variety over $F$, it suffices to prove that $X_{\overline{\eta}}$ is a primitive symplectic variety over $\overline{F}$. One way to proceed is to exploit the fact that we are working over $\mathbb{C}$ and to apply [undefaax, Lemma 2.1]. We can then find a closed point $s\in S$ such that $X_{s}$ is a complex primitive symplectic as in the statement of this lemma, an (abstract) field isomorphism $\phi:\mathbb{C}\to\overline{F}$, and an isomorphism of schemes $\alpha:X_{\overline{\eta}}\to X_{s}$ over $\mathop{\rm Spec}\nolimits(\phi)$. It follows that
@@
is a projective primitive symplectic variety over $\overline{F}$ by Lemma 2.3, and $b_{2}(X_{\eta})\geq 5$. Moreover, the isomorphism $\alpha$ induces isomorphisms between the divisor class groups and the Picard groups of $X_{s}$ and $X_{\overline{\eta}}$, and the following diagram is commutative.
@@
It follows that $X_{s}$ is $\mathbb{Q}$-factorial if and only if $X_{\overline{\eta}}$ is $\mathbb{Q}$-factorial. Since $K_{X_{\overline{\eta}}}=\alpha^{*}K_{X_{s}}$ we also see that the singularities of $X_{s}$ are terminal if and only if those of $X_{\overline{\eta}}$ are also terminal. Therefore, by [undefp, Lemma 6.18] and [undefan, Proposition 2.15], the singularities of $X_{\eta}$ are $\mathbb{Q}$-factorial and terminal, and that completes the proof. ∎
We now state a weaker version of [undefr, Lemma 2.17] in the singular setting, which will be useful in what follows.
@@
Since the interiors of the cones are the same, we only have to check that the boundary of the left-hand cone is included in the right-hand cone. Let $\alpha\in\partial\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$ and, without loss of generality, suppose that $\alpha$ is integral. Since $\overline{\mathop{\rm BAmp}\nolimits(X)}=\mathop{\overline{\rm Mov}}\nolimits(X)$ by [undefas, Proposition 5.8], there exists a projective primitive symplectic variety $X^{\prime}$ with $\mathbb{Q}$-factorial and terminal singularities, and a birational map $\phi:X\dashrightarrow X^{\prime}$ such that $\phi_{*}\alpha$ is integral and nef. Thanks to our assumptions, $\phi_{*}\alpha$ is semi-ample on $X^{\prime}$, and is therefore effective. Since $\phi$ is an isomorphism in codimension one, $\alpha$ is effective and this concludes the proof. ∎
The following result is an adaptation of [undefah, Proposition 5.5] to the singular setting.
@@
Let $f:X\to S$ be a projective fibration between $\mathbb{Q}$-factorial varieties, and let $X_{s}$ denote a very general fibre of $f$. Let $\eta\in S$ be the generic point of $S$, and denote by $j:X_{\eta}\to X$ the inclusion. Then the following assertions hold.
@@
is ample on $X_{\eta}$. Moreover, by [undefad, Corollaire 9.6.4], for each such integer $m$, there exists an open subset $U_{m}\subseteq S$ such that $\left(D+m^{-1}H\right)|_{X_{s}}$ is ample on $X_{s}$ for all $s\in U_{m}$. By (4.1), these open subsets are non-empty. Set $U:=\cap_{m\geq 1}U_{m}$. By construction, $D|_{X_{s}}$ is nef for all $s\in U$, and our hypothesis implies that $D|_{X_{s}}$ is also effective for all $s\in U$. Hence $D\in\mathop{\rm Eff}\nolimits(X/S)$, and therefore $D_{\eta}\in\mathop{\rm Eff}\nolimits(X_{\eta})$.
(2) We first prove the inclusion $\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})\subset\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$. Let $D_{\eta}\in\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})$. By Lemma 4.11, there exists $D\in\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)$ such that $D|_{X_{\eta}}=D_{\eta}$. By [undefah, Lemma 3.4], we have $D\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$, thus $D_{\eta}\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$. For the reverse inclusion $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})\subset\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})$, let $D_{\eta}\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$ and assume, without loss of generality, that $D_{\eta}$ is rational. By Lemma 4.11, we may choose $D\in N^{1}(X/S)_{\mathbb{Q}}$ such that $D|_{X_{\eta}}=D_{\eta}$. Let $H$ be an $f$-ample divisor on $X$. For any integer $m>0$, the divisor
@@
is big on $X_{\eta}$, thus we may write $(D+m^{-1}H)|_{X_{\eta}}=A_{m,\eta}+E_{m,\eta}$ with $A_{m,\eta}\in\mathop{\rm Amp}\nolimits(X_{\eta})$ and $E_{m,\eta}\in\mathop{\rm Eff}\nolimits(X_{\eta})$. Choose rational classes $A_{m}\in N^{1}(X/S)_{\mathbb{Q}}$ and $E_{m}\in\mathop{\rm Eff}\nolimits(X/S)$ such that $(A_{m})|_{X_{\eta}}=A_{m,\eta}$ and $(E_{m})|_{X_{\eta}}=E_{m,\eta}$. As above, for each integer $m>0$, there exists a non-empty open subset $U_{m}\subseteq S$ such that $(D+m^{-1}H)|_{X_{s}}=(A_{m}+E_{m})|_{X_{s}}$ is big for all $s\in U_{m}$. Setting again $U:=\cap_{m\geq 1}U_{m}$, it follows that $D|_{X_{s}}$ is pseudoeffective for all $s\in U$. By [undefar, Theorem 3.2], we may write $D|_{X_{s}}=P_{s}+N_{s}$, with $P_{s}\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{s})$ and $N_{s}\in\mathop{\rm Eff}\nolimits(X_{s})$. Since $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{s})\subseteq\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{s})$ by Lemma 4.15, it follows that $D|_{X_{s}}$ is effective. Therefore, $D$ is $f$-effective, and hence $D_{\eta}$ is effective.
As for the equality $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)=\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)$, the inclusion $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)\subset\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)$ can be proved in the same way as above. The reverse inclusion follows from [undefah, Lemma 3.4], which concludes the proof. ∎
We can now prove our final theorem.
@@
By Lemma 4.14, the generic fibre $X_{\eta}$ is a projective primitive symplectic variety over $F=\mathbb{C}(S)$ with $b_{2}(X_{\eta})\geq 5$ such that $X_{\overline{\eta}}$ has $\mathbb{Q}$-factorial terminal singularities. Theorem A.(2) implies that $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$ admits a rational polyhedral fundamental domain under the action of $\mathop{\rm Bir}\nolimits(X_{\eta})$. Under our assumptions, we have $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})=\mathop{\overline{\rm Mov}}\nolimits^{e}(X_{\eta})$ by Proposition 4.16. As observed in [undefah], the proof of [undefat, Theorem 6.1] shows that the weak cone conjecture 4.10 for $\mathop{\overline{\rm Mov}}\nolimits^{+}(X_{\eta})$ implies the weak cone conjecture for $\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$. Finally, since we also have $\mathop{\overline{\rm Mov}}\nolimits^{e}(X/S)=\mathop{\overline{\rm Mov}}\nolimits^{+}(X/S)$ by Proposition 4.16, it follows from [undefah, Lemma 2.18] that the relative movable cone conjecture holds for $f:X\to S$. Finally, by [undefah, Proposition 4.3], there are only finitely many small $\mathbb{Q}$-factorial modifications of $X$ over $S$ up to isomorphism over the base, and the relative nef cone conjecture for each of them. This completes this proof. ∎